OpenStudy (anonymous):

(cot^-1)(cosecx + cotx) differentiate w.r.t x

sam (.sam.):

try u sub \[u=\text{csch}(\cot (x))\text{ and }v=\text{}[\cot^{-1} (x)]\]

OpenStudy (anonymous):

\[cosec x + \cot x = \frac{1+ \cos x}{sinx}\] \[ = \frac{2\cos^2(x/2)}{2\sin(x/2)\cos(x/2)}\] =cot (x/2) Therefore, Now you can try :)

OpenStudy (anonymous):

@rroat

OpenStudy (anonymous):

A plot of the problem expression is attached.